# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ Appendix 0.A Additional Content for Membership Problem
    If $R=\{r_{ssq}\}$, then relating two variables $x$ and $y$ in both directions, i.e., putting $(r_{ssq},x,y)\in r_{\alpha^{\prime}}$ as well as $(r_{ssq},y,x)\in r_{\alpha^{\prime}}$ results in some substitution $h\in H$ to be valid if and only if $h(x)=h(y)$, as otherwise one can’t be a subsequence of the other. That concludes this case.
  3. 3.
    If $R=\{r_{ab}\}$, then valid substitutions of related variables result in them to have equal length, as abelian equivalence implies that. Hence, this case follows as in the case of the equal length relation $r_{|w|}$.
  4. 4.
    If $R=\{r_{perm}\}$, then we also get that two related variables must be substituted with equal length. As all substitutions must result in unary words consisting only of the letter $1$, this case also follows directly as the case of the equal length relation.
  5. 5.
    Also in the case of $R=\{r_{rev}\}$ we can use the argument that only unary substitutions are relevant in this reduction. As the reversal of a unary word is again the same unary word, this case also follows immediately.
  6. 6.
    Finally, in the case of $R=\{r_{*}\}$, as we said before, we may use this relation in both directions for each pair of variables, resulting effectively in equality of variables as in the case of the subsequence relation $r_{ssq}$.
This concludes all cases. For larger alphabets, the same construction may be used by trivially extending the word and the pattern by terminal letters made of up all additional letters in the alphabet. ∎
Proof of Proposition˜2:
Proof
We show this, also, by reducing from the $3\mathtt{-SAT}$ problem. So, as before, let $\varphi$ be a $3\mathtt{-SAT}$ formula and let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$. Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$. For each clause $C_{j}\in C$ for $j\in[n]$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$. W.l.o.g. assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause). Define $n+1$ functions $f_{j^{\prime}}$ with $j_{j^{\prime}}\in[n]_{0}$ such that $f_{j^{\prime}}:\mathcal{X}\rightarrow X$ with $f_{j^{\prime}}(X_{i})=u_{i,j^{\prime}}$ and $f_{j^{\prime}}(\overline{X_{i}})=v_{i,j^{\prime}}$. Assume $R=\{r\}$ for $r\in\{r_{com^{+}},r_{com^{*}}\}$. Construct the word
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
with
$s_{i}$ $=1\#1\text{, and}$
$t_{j}$ $=1^{10}\#1^{10}\#1^{10}$
for $i\in[m]$ and $j\in[n]$ as well as the relational pattern $(\alpha,r_{\alpha})$ such that $\alpha$
$\alpha=\#\#\ \alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
for $i\in[m]$ and $j\in[n]$ with
$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}}),\text{ and}$
$\alpha_{2,j}$ $=z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{j,3}\ f_{j}(X_{j,3})\ z_{j,4},$
and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$. Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp. $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$. Also notice, that each variable $z_{i,j}$ is independent and may be substituted to anything.
Now, assume there exists an assignment of the variables in $\mathcal{X}$ such that $\varphi$ is satisfied. We will now construct a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$. For each $i\in[m]$, if $X_{i}$ is true (hence, $\overline{X_{i}}$ is false), set $h(u_{i,j})=1$ and $h(v_{i,j})=\#1$, otherwise set $h(u_{i,j})=1\#$ and $h(v_{i,j}=1)$ for $j\in[n+1]$. This results in $h(\alpha_{1,i})=1\#1$ for all $i\in[m]$. As there is at least one variable assigned to true for each clause $j\in[n]$, we know that $|h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))|_{\#}\leq 2$. As there is at most two $1$’s between two occurrences of #\# and at most $3$ occurrences of $1$’s before or after the first or last occurrence of #\# in $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))$, we always find a non-empty substitution of all variables $z_{j,1}$ to $z_{j,4}$ such that $h(\alpha_{2,j})=\#\#\ 1^{10}\#1^{10}\#1^{10}$. All substitutions done by $h$ result in related variables to be substituted equally, hence, this satisfies the relation $r$, and by that there exists some $r_{\alpha}$ valid $h\in H$ such that $h(\alpha)=w$.
Now assume $w\in L_{NE}(\alpha,r_{\alpha})$. Then there exists some $r_{\alpha}$ valid substitution $h\in H$ such that $h(\alpha)=w$. We know that each occurrence of the factor ##\#\# in $w$ must be mapped by one occurrence of the factor ##\#\# in $h(\alpha)$. As the same number of ##\#\# occur as terminal letters in $\alpha$ as there are occurrences of it in $w$, we know that these occurrences must align. Hence, we get that for all $i\in[m]$ and $j\in[n]$ we have
$h(\alpha_{1,i})=h(u_{i,1}v_{i,1})=1\#1=s_{i}$

### v2

@@ 4.2 The Erasing Cases
    If $|\Sigma|=1$, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
For the undecidability proofs, we use the same framework as in the result from Freydenberger and Reidenbach [13] that shows undecidability in the erasing case under $r_{=}$, for all alphabets $\Sigma$ with $|\Sigma|\geq 2$, as a starting point, but we need to adapt it at several points in a more involved way to get the following theorem.
Theorem 4.3(✂0.B.3).
Let $\bar{r}\in\{r_{rev},r_{ab}\}$ and let $R=\{\bar{r}\}$. For all alphabets $\Sigma$ with $|\Sigma|\geq 2$, given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is generally undecidable whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
Analogously to the proof of Theorem˜4.1, a full proof sketch for Theorem˜4.3 is given in Section˜0.B.3.
5 The Membership Problem
In this final section, we briefly discuss the membership problem – also known as the matching problem – for relational pattern languages for all relations in $\mathcal{R}$ and general patterns. For the case of $r_{=}$, NP-completeness of the membership was originally shown in [3] and [22] for the non-erasing and erasing cases respectively. In fact, independent communities have shown this under several considered restrictions (cf. [11, 28]). As shown in this section, all but one cases remain NP-complete.
Proposition 2.
Let $r\in\mathcal{R}$ and $R=\{r\}$. Let $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. Deciding $w\in L_{X}(\alpha,r_{\alpha})$, for $X\in\{E,NE\}$, is in NP.
Indeed, NP-containment follows by the known arguments from the cases arising in $r_{=}$. We can guess a substitution $h$ with $|h(\alpha)|=|w|$, check the relational constraints in polynomial time, and finally check whether $h(\alpha)=w$. Regarding NP-hardness, we revisit the exposition given in [28], where a reduction is given from the known NP-complete positive 1-in-3-SAT (no negated variables, 3-CNF clauses, satisfiable with exactly one literal per clause set to true) problem. For example, assuming $\Sigma=\{0,\#\}$, the formula
$\{\{X_{1},X_{2},X_{3}\},\{X_{2},X_{4},X_{5}\},\{X_{1},X_{4},X_{5}\},\{X_{3},X_{4},X_{6}\}\}$
is in 3-CNF and 1-in-3 satisfiable if and only if the word w=0#0#0#0w=0\#0\#0\#0 is in the erasing pattern language of the pattern $\alpha=x_{1}x_{2}x_{3}\#x_{2}x_{4}x_{5}\#x_{1}x_{4}x_{5}\#x_{3}x_{4}x_{6}$. Using $w=0^{4}\#0^{4}\#0^{4}\#0^{4}$, this construction works for the NE case as well. For most of the relations considered in this paper, this framework suffices.
Lemma 5(✂0.C.1).
Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*},r_{com^{+}},r_{\varepsilon_{=}}\}$ s.t. $R=\{r\}$ and let $|\Sigma|\geq 2$. For $w\in\Sigma^{*}$, $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding $w\in L_{E}(\alpha,r_{\alpha})$ is NP-hard. If $r\notin\{r_{com^{+}},r_{\varepsilon_{=}}\}$, deciding $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard.
This leaves us with the cases of $r_{com^{+}}(r_{com^{*}})$ in the non-erasing case and with $r_{com^{*}}$ in the erasing case. For these, a slightly different reduction is necessary but possible to achieve. In Appendix 0.C.2, a reduction from the 3-SAT problem is given, concluding the following result.
Theorem 5.1(✂0.C.2).
Let $r\in\mathcal{R}$ s.t. $R=\{r\}$ and assume $|\Sigma|\geq 2$. For $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding $w\in L_{E}(\alpha,r_{\alpha})$ is NP-complete. If $r\neq r_{\varepsilon_{=}}$, deciding $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-complete.
Finally, we recall from Remark˜1 that the relation $r_{\varepsilon_{=}}$ has no effect in the non-erasing case, resulting in each pattern language $L_{NE}(\alpha,r_{\alpha})$, for $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $R=\{r_{\varepsilon_{=}}\}$, being equal to the pattern language $L_{NE}(\alpha)$ of the regular pattern $\alpha$. The membership problem for regular patterns has been shown to be decidable in linear time by Manea et al. in [27]. Hence, the following immediately holds.
Corollary 3.
Let $R=\{r_{\varepsilon_{=}}\}$. For $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we can decide in linear time whether $w\in L_{NE}(\alpha,r_{\alpha})$.
6 Final Discussion
We considered the three main decision problems equivalence, inclusion, and membership problems for relational pattern languages under various typical binary relations, i.e., $r_{=}$, $r_{ssq}$, $r_{ab}$, $_{perm}, rr$$_{rev}$, $r_{com^{+}}$, $_{com^{*}}, rr_{*}$, and $r_{\varepsilon_{=}}$. While obtaining a significant number of results, different research directions building on the established foundation can be worthwhile pursuing.
Regarding the equivalence problem, we obtained decidability in polynomial time for $r_{\varepsilon_{=}}$, $r_{ab}$ and $r_{com^{+}}$ in the non-erasing case, similar to the known cases $r_{=}$ and $r_{|w|}$. For the relations $r_{rev}$, $r_{ssq}$, and $r_{*}$ we obtained strict structural implications of patterns with equal languages. A clear answer regarding (efficient) decidability (except the known general decidabilit for $r_{rev}$), however, probably requires a more in depth analysis. For the erasing cases, most significantly, the relation $r_{\varepsilon_{=}}$ results in decidability. Together with the known result for $r_{|w|}$, this marks one of the few relations with a clear positive answer with respect to decidability in the erasing case. We recall that this problem for $r_{=}$ remains the main open problem for decision problems in pattern languages. In addition to that, we could establish that this problem under the relations $_{rev}, rr_{*}$, and $r_{ssq}$ must be at least as hard to solve as for $r_{=}$, extending the decidability question in that direction.
For the inclusion problem, we showed that $r_{\varepsilon_{=}}$ results in decidability. Together with $r_{|w|}$, this is the only other relation currently known to do so. In contrast, inclusion is undecidable for the relational patterns under $r_{=}$, $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$. These undecidability results arise either because the relations can simulate $r_{=}$ in general, as with $r_{ssq}$ and $r_{*}$, or because they still allow for the key properties used in established reductions to be enforced, such as equality of letters or unary words, as with $r_{ab}$ and $r_{rev}$. It remains open whether similar results can be obtained for structurally weaker relations. For the remaining cases $r_{perm}$, $r_{com^{*}}$, and $r_{com^{+}}$, the crucial reduction properties do not seem to be achievable in an obvious way. The relation $r_{perm}$ does not enforce equal letters or equal unary words over alphabets of size at least two, while $r_{com^{+}}/r_{com^{*}}$ do not even enforce equal lengths of substituted words. Hence, the boundary between (un)-decidability for weaker relations remains an interesting open question.
Finally, for the membership problem, we observed that in all considered cases (except $r_{\varepsilon_{=}}$ in the non-erasing case), that is, for all relations in $\mathcal{R}$ in the erasing case and all relations in $\mathcal{R}\setminus\{r_{\varepsilon_{=}}\}$ in the non-erasing case, the problem turns out to be NP-complete. Interestingly, this even holds for structurally very weak relations such as $r_{\varepsilon_{=}}$ or $r_{com^{*}}/r_{com^{+}}$. As generally assumed, searching for words in sets of texts that are compactly described by, e.g., a relational pattern of almost any kind, turns out to be a generally computationally hard problem.
This concludes this paper. Again, for a visual summary, please see Table 1.
Acknowledgements.
We thank the anonymous reviewers of the previous version of this paper for their valuable comments and suggestions. In particular, their remarks on the membership problem helped us to better position our results within the context of more recent research.
Appendix 0.A Additional Content for the Equivalence Problem
0.A.1 Proof of Lemma˜2
Proof.
($\Rightarrow$): First, let $w\in L_{NE}(\alpha,r_{\alpha})$. Then there exists some $r_{\alpha}$-valid substitution $h$ with $h(\alpha)=w$. Let $(r,x_{i},x_{j})\in r_{\alpha}$, for two variables $x_{i},x_{j}\in\operatorname{var}(\alpha)$. Then $(h(x_{i}),h(x_{j}))\in r$ by definition. As $r$ is an equivalence relation. we also have $(h(x_{j}),h(x_{i}))\in r$, $(h(x_{i}),h(x_{i}))\in r$, and, if there exists some $(r,x_{j},x_{k})\in r_{\alpha}$ (hence $(h(x_{j}),h(x_{k}))\in r$), we also have $(h(x_{i}),h(x_{k}))\in r$. Hence, $h$ satisfies all constraints in $r_{\alpha}]$ and thereby is $r_{\alpha}]-$valid, resulting in $w\in L_{NE}(\alpha,[r_{\alpha}])$.
($\Leftarrow$): For the other direction, let $w\in L_{NE}(\alpha,[r_{\alpha}])$ and let $h$ be an $r_{\alpha}]-$valid substitution such that $h(\alpha)=w$. Notice that $r_{\alpha}\subseteq[r_{\alpha}]$. Hence, all constraints in $r_{\alpha}$ are satisfied by $h$, resulting in $h$ being $r_{\alpha}$-valid. Thus, $w\in L_{NE}(\alpha,r_{\alpha})$. ∎
0.A.2 Proof of Theorem˜3.1
Proof.
We obtain the first direction immediately from Lemma˜1. For the other direction, assume $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. By Lemma˜2 we know $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$ and $L_{NE}(\beta,r_{\beta})=L_{NE}(\beta,[r_{\beta}])$. As $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$, we trivially have $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$. Thus, $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$. ∎
0.A.3 Proof of Corollary˜1
Proof.
Since every relation in $r_{=},r_{ab},r_{con^{+}}$} is an equivalence relation enforcing equality of related letters, we first normalize $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ by renaming variables in order of first occurrence to $x_{1},x_{2},\dots$. We then test whether $\alpha=\beta$. If not, we return false. Otherwise, we compute $r_{\alpha}]$ and $r_{\beta}]$ via their reflexive, symmetric, and transitive closures by a single scan of the patterns. These classes have total size at most the number of variables, so this also takes linear time. We return true exactly if $r_{\alpha}]=[r_{\beta}]$, and false otherwise, by Theorem˜3.1. ∎
0.A.4 Proof of Theorem˜3.2
Proof.
Again, consider $(\alpha,r_{\alpha})$ and for each subset $S\subseteq\mathcal{C}_{\alpha}$, replace the variables $v\in\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by any word in $\Sigma^{+}$ and any other variable $v\notin\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by $\varepsilon$. We call the resulting language $L_{E}(\alpha,r_{\alpha},S)$. Using that, we obtain $L_{E}(\alpha,r_{\alpha})=\bigcup_{S\subseteq\mathcal{C}_{\alpha}}L_{E}(\alpha,r_{\alpha},S)$. Do the same for $\beta$. Note there exist at most $2^{k_{\alpha}}\leq 2^{|\alpha|}$ subsets $S$ for $\alpha$ and at most $2^{k_{\beta}}\leq 2^{|\beta|}$ subsets for $\beta$.
Here, we provide a more detailed construction of the previously mentioned NFAs: Starting with $(\alpha,r_{\alpha})$, for each such $S$ defined above, we can take a sub-pattern $\alpha_{S}$ of $\alpha$ that only contains the variables that are not substituted by the empty word. As $r_{\varepsilon_{=}}$ imposes no structural restriction other than being substituted (non)emptily commonly, essentially $\alpha_{S}$ is a regular pattern and we obtain the relationship $L_{NE}(\alpha_{S})=L_{NE}(\alpha,r_{\alpha},S)$. So, for each such regular sub-pattern $\alpha_{S}$, we can create an NFA $A_{\alpha_{S}}=(Q,\Sigma,\delta,q_{0}),F$ with $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})$ using a straightforward construction that represents that regular pattern. Consider the following sketch. We set $Q=\{q_{\alpha_{S},0},...,q_{\alpha_{S},|\alpha_{S}|}\}$ where each state $q_{\alpha_{S},i}$ represents a position between some letters in the pattern $\alpha_{S}$ (or a position in front of the first letter or after the last letter). Then we set the state corresponding to the last position in $\alpha_{S}$ (i.e., $q_{\alpha_{S},|\alpha_{S}|}$) to the only final state. Hence, $F=\{q_{\alpha_{S},|\alpha_{S}|}\}$. For all other positions $i<|\alpha_{S}|$, if $\alpha_{S}[i+1]$ is some terminal letter $\mathtt{a}\in\Sigma$, we add only one outgoing transition and set $\delta(q_{\alpha_{S},i},\mathtt{a})=\{q_{\alpha_{S},i+1}\}$. On the other hand, if $\alpha_{S}[i+1]\in\operatorname{var}(\alpha_{S})$ is some variable, add two outgoing transitions for each letter, one going to the state representing the next position, the other one staying in the same state. Thus, set $\delta(q_{\alpha_{S},i},\mathtt{a})=\{q_{\alpha_{S},i},q_{\alpha_{S},i+1}\}$, for all $\mathtt{a}\in\Sigma$. As $\alpha_{S}$ is a regular pattern as established before, the relationship $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})$ is apparent. Now, for each $S\subseteq\mathcal{C}_{\alpha}$, we can construct such an NFA $A_{\alpha_{S}}$. Finally, we can combine all of these NFA into one common NFA $A_{\alpha}$ where a new initial state is introduced that is connected to each initial state of the sub-automata $A_{\alpha_{S}}$ by an $\varepsilon$-transition, and each final state is preserved. So, using one outgoing transition of the initial state of $A_{\alpha}$, we choose which blocks of variables $S$ are to be substituted emptily together and which are not. Then, the respective sub-automaton $A_{\alpha_{S}}$ produces some word of the corresponding sub-language $L_{E}{\alpha,r_{\alpha_{S}}}$. In total, we see that the NFA $A_{\alpha}$ can be constructed and represents $L(A_{\alpha})=L_{E}{\alpha,r_{\alpha}}$. So the same for $(\beta,r_{\beta})$ by constructing an NFA $A_{\beta}$ analogously. It is well known that checking equivalence of the languages of two NFA is PSPACE-complete (see, e.g., [1]). Hence, we can check $L(A_{\alpha})=L(A_{\beta})$ and use the result to decide $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$. ∎
0.A.5 Proof of Lemma˜3
Proof.
(1): Let $\alpha=x_{1}x_{2}\dots x_{n_{\alpha}}$ and $\beta=y_{1}y_{2},\dots y_{n_{\beta}}$ be terminal-free and let $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\beta}$ be the equivalence classes. Note that the empty word can be obtained by both patterns. Thus, in the following, we focus on non-empty words.
Consider the pattern $\alpha$. Let $S\subseteq\mathcal{C}_{\alpha}$ with $S\neq\emptyset$ be the subset of equivalence classes chosen to be non-empty. Assume that all variables in the remaining classes are replaced by $\varepsilon$. Since each non-empty variable contributes at least one letter to the resulting word but has no other restriction, we can build all words of length at least $\sum_{C_{\alpha,i}\in S}|C_{\alpha,i}|$. Now, when taking all possible subsets $S$ into account, no word $w$ of positive length below $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|$ is possible. Thus, the language of accepted words by $\alpha$ is
$L_{E}(\alpha,r_{\alpha})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}.$
Respectively, the language of accepted words by $\beta$ is
$L_{E}(\beta,r_{\beta})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
Now, $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|=\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$ if and only if both patterns are equivalent. This property can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
(2): Assume $|\Sigma|=1$ and $a\in\Sigma$. Note that the patterns $\alpha$ and $\beta$ must have the same number of terminals in order to be equivalent. Otherwise we could set all variables to $\varepsilon$ and obtain distinct words. Let $\ell$ be the number of terminals in each pattern. By the same arguments as in (1), either all variables are empty, i.e., $\{a^{\ell}\}\in L_{E}(\alpha,r_{\alpha})$ and $\{a^{\ell}\}\in L_{E}(\beta,r_{\beta})$, or a non-empty subset $S$ of equivalence classes contributes $\sum_{C_{\alpha,i}\in S}|C_{\alpha,i}|$ (respectively $\sum_{C_{\beta,i}\in S}|C_{\beta,i}|$) $a$’s to the words. As we consider a single-letter alphabet, the language of accepted words by $\alpha$ is
$L_{E}(\alpha,r_{\alpha})=\{a^{\ell}\}\cup\{w\mid|w|\geq\ell+\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}.$
Respectively, the language of accepted words by $\beta$ is
$L_{E}(\beta,r_{\beta})=\{a^{\ell}\}\cup\{w\mid|w|\geq\ell+\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
We obtain that both patterns are equivalent if and only if they have the same number of terminals and $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|=\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$. This can again be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
(3): Assume there exists a letter $c\in\Sigma$ that does not appear as a terminal in either $\alpha$ or $\beta$. If all variables are set to $\varepsilon$, both patterns generate a word that matches their terminals. Hence, a necessary condition for equivalence is that the terminal words of $\alpha$ and $\beta$ are identical. This, we assume in the following and focus on words where at least one variable is non-empty.
Let $m$ be the number of terminals in $\alpha$ or $\beta$. We write
$\alpha=V_{\alpha,0}T_{1}V_{\alpha,1}T_{2}\dots T_{m}V_{\alpha,m},$
where $T_{j},j\in\{1,\dots,m\}$ are blocks terminals and $V_{\alpha,j},j\in\{0,\dots,m\}$ are blocks of variables. Since the terminals of both patterns are identical, $\beta$ has the same decomposition but the blocks of variables may be different, i.e.,
$\beta=V_{\beta,0}T_{1}V_{\beta,1}T_{2}\dots T_{m}V_{\beta,m}.$
First, consider the pattern $\alpha$. For each equivalence class $C_{\alpha,i}$, let $\mathbf{occ}_{\alpha,i}(j)$ be the number of occurrences of that class in block $V_{\alpha,j}$ and define
$\mathbf{occ}_{\alpha,i}:=(\mathbf{occ}_{\alpha,i}(0),\mathbf{occ}_{\alpha,i}(1),\dots,\mathbf{occ}_{\alpha,i}(m)).$
If a subset $S\subseteq\mathcal{C}_{\alpha}$ is non-empty, the minimal number of letters in block $V_{j}$ is $\sum_{C_{\alpha,i}\in S}\mathbf{occ}_{\alpha,i}(j)$. Thus, every word of length at least $\sum_{C_{\alpha,i}\in S}\mathbf{occ}_{\alpha,i}(j)$ can be generated in that block.
We say that an equivalence class $C_{\alpha,i}$ is dominated if there exists a set $S^{\prime}$ of other classes, i.e., $S^{\prime}\subset\mathcal{C}_{\alpha}$ and $C_{\alpha,i}\notin S^{\prime}$, such that for all blocks $j\in\{0,\dots,m\}$ we have
$\mathbf{occ}_{\alpha,i}(j)=0\Longleftrightarrow\sum_{C_{\alpha,i^{\prime}}\in S^{\prime}}\mathbf{occ}_{\alpha,i^{\prime}}(j)=0$